Some towns and cities have passed statutes requiring higher wages. This homework assignment examines the economic resuts of these statutes. Some argue that the new laws are aimed atlarge firms thatemploy manyworkers and earn largeprofits from each worker. Consider the scenario below, which reflects the more typical case. We do not judge the morality of these statutes or whether low paid workers are exploited by firms. Exploit is a normative tem imply ngthat firms un ary appropriate tewealth of therworkers.Landsburg does not deal with normative issues and does not use this term. Scenao#2 Afr an hour, so its net profit is $10.00 an hour. A new statute sets a minimum wage of $10.00 an hour for firms with 30 or more employees. The advocates of the new statute say the firm will now pay $10.00 an hour, so its net proft is $6.00 an hour and te workers receive a better wage. with 100 workers no pay them $6.00an hour. From each worker, thef gains$16.00 Scenario #1 : A fmm with 30 workers now pays tem $0.00 an hour From each worker, the frm gains S 11.00 an hour, so its net profit is $5.00an hour. A new statute sets a minimum wageof$10.00an hour for firms with 30 or more employees. The advocates of the new statute say the firm will now pay $10.00 an hour, so its net profit is $1.00 an hour and its workers receive a better wage. If we examine the cases outined above, the fim earns greater profits byretaining its workersandpayingthem $10.00 an hour than bylaying off 75 workers and paying the rest of the workers $6.00 an hour We examine three cases: a: The firm pays $6.00 an hour to thirty workers. B: The fim pays $10.00 an hour to thirty workers. y: The firm lays off 5 workers and pay $6.00 an hour to 25 workers. To examine this scenario, we assume the firm would pay $6.00 an hour if it moved its operations toa neighboring town (Town zz), but the move would cost $250,000. Assume the revenue of the fim does not change tmoves to ZZ. A. Whatis the dfference in labor costs if the fim moves to ZZ?(S10.00-$6.00)x 100workers each hour.) B. What is the difference in net profit? C. Assuming the work week is 40 hours, what is the difference in weekly profits? D. How many weeks would it take the firm to recoup the moving costs? E. What is the effect of the new statute on unemployment? For each case, work out the Total revenue of the firm Total labor costs of the firm. Net profit of the firm. For case α, the total revenue is $11 × 30 = $330 an hour, te total labor costs are $6 30 = $180; and the net profit is $330 $180 $150. For the homework assignment, determine the corresponding figures for cases and y and answer the tems below A. Total revenue when it pays $6.00 an hour to 30 workers? (30x $11.00) B. Total labor cost when it pays $6.00 an hour to 30 workers? (30x $6.00) C. Total profit when it pays $6.00 an hour to 30 workers? (30x $5.00) D. Total labor cost if it pays $10.00 an hour to 30 workers? (30 x $10.00) E. Total profit if t pays $10.00 an hour to 30 workers? (30x $1.00) F. Total revenue f it lays off 5 workers and pays $6.00 to 25 workers? (25 $11.00) G. Total abor cost if it lays off 5 workers and pays $6.00 to 25 workers? (25x $6.00) H. Total profit if lays off 5 workers and pays $6.00 to 25 workers? (25 $5.00) I. What is the expected response of a profit maximizing firm to the new labor law? J. What is the expected change to unemployment? K. Do workers gain from the new law? L. Do firms gain from the new law?
Answer #1

Alpha is already given

A. TR, $11 * 30 $330

B. TC- $6 * 30 $180

large Profit, pi_{alpha }= TR_{alpha } - TC_{alpha}

C. Profit (alpha) = $ 330 - 180 = $ 150

Beta

A. TR3 = $11 * 30-$330

D. TC3 = $10 * 30-$300

Profit, T3TR-TCs

E. Profit(beta) = $ 330 - 300 = $ 30

Gamma: 25 workers and wage $6

F. TR$11 25$275

G. TC, $6 * 25-$ 150

Prof it.т, = TR.-TC.

H. Profit (gamma) = $ 275 - 150 = $ 125

I. Expected response of the profit maximizing firm to the new law.

The firm will layoff 5 workers and continue to pay $6 per worker to the remaining 25 workers.

J. Expected change in unemployment

5 workers will become unemployed.

K. Do workers gain from the new law?

In case of Beta they gain.

In case of Gamma no.

L. Do firms gain from the new law?

In case of Beta - No

In case of Gamma- No (Since it is less than alpha).

Case 2.

Present scenario

TR = $16*100 = $ 1600

TC = $6*100 = $ 600

Profit = 1600 - 600 = $ 1000

When new law is imposed

TR' = $ 16*100 = $ 1600

TC' = $ 10*100 = $ 1000

New Profit = $ 600

When they layoff 75 workers and continue to pay $6 to remaining 25 workers

TR" = $16*25 = $ 400

TC" = $ 6*25 = $ 150

PROFIT" = $ 250

Now when the firm moves to town ZZ

A. Difference in Labour cost = ($10-6)*100 = $ 400

B. Difference in net profit

Profit before moving out of town = $ 600

Profit after moving out of town = $ 1000

Difference in profit = $ 1000 - 600 = $ 400

C. Week = 40 hours

1 hour profit (before) = $ 600

Weekly profit (before) = 40*600 = $ 24,000

1 hour profit (after) = $ 1000

Weekly profit(after) = 40*1000 = $ 40,000

Difference in weekly profit = $ 16,000

D. Payback period

$250,000 $16000 Numberofweeks

Number of weeks to recoup = 15.625 weeks

E. Due to new rule. The firms will move out of the town in order to earn a higher profit. If the firms continue to be in the same town then they won't layoff workers but their net profit will decline.

Please contact if having any query thank you.